OpenStudy (anonymous):

Could you solve 175r^5/sqrt(64^17) and show all your work.

OpenStudy (abb0t):

Is that a question or a challenge?

OpenStudy (anonymous):

both I guess

OpenStudy (anonymous):

please help

OpenStudy (jhannybean):

\[\frac{ 175r^5 }{ \sqrt{64^{17}} }\] right?

OpenStudy (jhannybean):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{ 175r^5 }{ \sqrt{64^{17}} }\] \[\large \frac{ 175r^5 }{ \left(\sqrt{64}\right)^{17} }\] \[\large \frac{ 175r^5 }{ 8^{17} }\] you can rewrite \(\large 8^{17}\) as \(\large (2^3)^{17} = 2^{51}\), but it's pretty much simplified at this point